# Phase structure of (3+1)-dimensional dense two-color QCD at $T=0$ in the strong coupling limit with tensor renormalization group

### v1

@@ Abstract
We investigate the phase structure of the (3+1)-dimensional strong coupling two-color QCD at zero temperature with finite chemical potential using the tensor renormalization group method. The chiral and diquark condensates and the quark number density are evaluated as a function of the chemical potential. Our results are compared with the previous analytical results using the mean field approximation. The critical exponents associated with the diquark condensation are also discussed.
preprint: UTHEP-811, UTCCS-P-169
1 Introduction
@@
The complex action problem with finite chemical potential $\mu$ does not allow the standard Monte Carlo simulation to investigate the phase structure of QCD in the finite baryon density. On the other hand, the simulation of two-color QCD (QC2D) with $\mu\neq 0$ is free from the complex action problem, so that it provides us an opportunity to understand the physics of matter under the finite density condition. Furthermore, in the strong coupling limit, we can use analytical methods to investigate the phase structure. In the 1980s, there were several mean field (MF) studies on chiral and diquark condensates in the strong coupling limit of the finite density QC2D with staggered quarks Dagotto:1986gw ; Dagotto:1986xt ; Klatke:1989xy . The results were later refined by Nishida, including the QCD case Nishida:2003uj ; Nishida:2003fb .
The tensor renormalization group (TRG) method  [^1] In this paper, the “TRG method” or the “TRG approach” refers to not only the original numerical algorithm proposed by Levin and Nave Levin:2006jai but also its extensions PhysRevB.86.045139 ; Shimizu:2014uva ; PhysRevLett.115.180405 ; Sakai:2017jwp ; PhysRevLett.118.110504 ; Hauru:2017tne ; Adachi:2019paf ; Kadoh:2019kqk ; Akiyama:2020soe ; PhysRevB.105.L060402 ; Akiyama:2022pse . may be qualified to investigate the phase structure of finite density QCD: This method, which is in principle free from the sign problem or the complex action problem, has been successfully applied to various types of models with these problems Shimizu:2014uva ; Shimizu:2014fsa ; Kawauchi:2016xng ; Kawauchi:2017dnj ; Yang:2015rra ; Shimizu:2017onf ; Takeda:2014vwa ; Kadoh:2018hqq ; Kadoh:2019ube ; Takeda:2019idb ; Kuramashi:2019cgs ; Akiyama:2020ntf ; Akiyama:2020soe ; Nakayama:2021iyp ; Bloch:2021mjw ; Bloch:2022vqz ; Akiyama:2023hvt ; Akiyama:2024qer ; Luo:2022eje ; Hite:2024ulb ; Luo:2024lbh ; Luo:2025qtv . As a preparatory study before exploring (3+1)-dimensional ((3+1)$d$) finite density QCD, it would be a good idea to apply the TRG method to the strong coupling limit of (3+1)$d$ finite density QC2D, which becomes much simpler with vanishing gauge action. Firstly, we already have some insight into the phase structure thanks to a recent study of (1+1)$d$ finite density QC2D with the TRG method, including the finite coupling case Pai:2024tip . Secondly, it is useful to compare the TRG results with the MF ones Nishida:2003uj . Especially, the TRG method is allowed to investigate the phase structure at zero temperature, where the standard Monte Carlo approach is hardly accessible. Thirdly, the computational techniques employed in our previous study on the (3+1)$d$ finite density Nambu-Jona-Lasinio model Akiyama:2020soe are readily applicable to the strong coupling case.
In this paper, we investigate the phase structure of the strong coupling limit of the finite density QC2D with the staggered quark, measuring the chiral and diquark condensates and the quark number density as a function of $\mu$. Our results are compared with the analytical results obtained by the MF method and the $1/d_{s}$ expansion with $d_{s}$ the spatial dimension Nishida:2003uj .
@@
This paper is organized as follows. In Sec. 2, we define the action of the QC2D with finite chemical potential and give the tensor network representation. In Sec. 3 we measure the chiral and diquark condensates and the quark number density as a function of the chemical potential. The critical exponents are also determined. The results are compared with the analytical ones in Ref. Nishida:2003uj . Section 4 is devoted to summary and outlook.
2 Formulation and numerical algorithm
2.1 Strong coupling QC2D
@@
We consider the finite density QC2D on a (3+1)$d$ lattice $\Lambda_{3+1}=\{(n_{1},n_{2},n_{3},n_{4})\ |n_{1,2,3}=1,\dots,N_{s}\,,n_{4}=1,\dots,N_{\tau}\}$ whose volume is $V=N_{s}^{3}\times N_{\tau}$. The lattice spacing $a$ is set to $a=1$ unless necessary. The action consists of the gluonic part and the fermionic one. For the former, the simple plaquette action is given by
@@
with $m$ the quark mass. $\chi=(\chi^{1},\chi^{2},\dots,\chi^{N_{c}})$ is a $N_{c}$ component Grassmann variables, and $\eta_{\nu}(n)$ is the staggered sign function defined by $\eta_{\nu}(n)=(-1)^{n_{1}+\cdots+n_{\nu-1}}$ with $\eta_{1}(n)=1$. Note that Eq. (2) is invariant under the following continuous chiral transformation at $m=\mu=0$:
$\chi(n)$ $\rightarrow$ $e^{i\alpha\epsilon(n)}\chi(n),$ (3)
$\bar{\chi}(n)$ $\rightarrow$ $\bar{\chi}(n)e^{i\alpha\epsilon(n)}$ (4)
with $\alpha\in\mathbb{R}$ and $\epsilon(n)=(-1)^{n_{1}+n_{2}+n_{3}+n_{4}}$.
@@
where $S_{D}$ is a source term to detect the diquark condensate given by
$S_{D}=\frac{\lambda}{2}\{\chi^{t}\tau_{2}\chi+\bar{\chi}\tau_{2}{\bar{\chi}}^{t}\}$ (6)
@@
2.2 Grassman tensor network representation
In order to apply tensor network methods to this system, it is necessary to rewrite the partition function $Z$ in Eq. (5) as a product of local tensors. Since the theory involves fermionic degrees of freedom, we must properly incorporate their anticommuting nature. To this end, we employ a Grassmann tensor network formulation Akiyama:2020sfo , in which the local tensors are defined using auxiliary Grassmann variables.
For clarity, let us consider the case $\lambda=0$, which simplifies the discussion. The extension to $\lambda\neq 0$ is straightforward. To obtain the Grassmann tensor network representation of the partition function in Eq. (5), We introduce two-component auxiliary Grassmann variables $\zeta_{\nu}$ and $\xi_{\nu}$ $(\nu=1,2,3,4)$ Akiyama:2020sfo . As a basic tool, for any $N$-component Grassmann variable $\theta_{i}$, one can use the identity
@@ 2.2 Grassman tensor network representation
In the above, we have introduced the fermion occupation numbers $i^{\prime 1}_{\nu},i^{\prime 2}_{\nu},j^{\prime 2}_{\nu},j^{\prime 1}_{\nu},j^{1}_{\nu},j^{2}_{\nu},i^{1}_{\nu},i^{2}_{\nu}\in\{0,1\}$ $(\nu=1,2,3,4)$, which arise from the Taylor expansion of Eq. (10), and the auxiliary integer variables $a_{\nu},b_{\nu},c_{\nu},d_{\nu}\in\{1,2\}$ to sum out the color indices of the staggered fields $\chi$ connected to the link variables $U_{\nu}$. The summation $\sum_{\{i,j\}}$ denotes $\left(\prod_{\nu}\sum_{i^{\prime 1}_{\nu},i^{\prime 2}_{\nu},j^{\prime 1}_{\nu},j^{\prime 2}_{\nu},j^{1}_{\nu},j^{2}_{\nu},i^{1}_{\nu},i^{2}_{\nu}}\right)$. Here we omit the explicit spacetime dependence except for $\eta_{\nu}(n)$. This implies that there are eight types of initial tensors, depending on the staggered sign Akiyama:2020soe .
@@
where $F^{a_{\nu}b_{\nu}c_{\nu}d_{\nu}}_{i^{\prime 2}_{\nu}i^{\prime 1}_{\nu}j^{\prime 2}_{\nu}j^{\prime 1}_{\nu}}$ indicates the results of the integration using Weingarten calculus Weingarten:1977ya . After integrating out the gauge fields, all that remains is to integrate out $\chi$ and $\bar{\chi}$. A straightforward calculation yields the final form of the Grassmann tensor $\mathcal{T}$:
@@
The tensor $\tilde{F}$ is a redefined version of $F$ that incorporates $a,b,c,d$ dependence of the sign factors arising from the reordering of $\chi^{2},\chi^{1},\bar{\chi}^{2},\bar{\chi}^{1}$ in the integration procedure. The tensor $R$ likewise includes sign factors originating from the rearrangement of the auxiliary Grassmann fields $\zeta,\xi,\bar{\zeta},\bar{\xi}$ to the form of Eq. (14). The explicit definitions of $F$, $\tilde{F}$, and $R$ are provided in the Appendix. Using these definitions, the partition function $Z$ can be expressed as a Grassmann tensor network Kadoh:2018hqq ,
@@
To evaluate Eq. (16), we have employed the Grassmann-anisotropic TRG method Adachi:2019paf ; Akiyama:2020soe with the bond dimension $D$ incorporating the multi-GPU parallelization technique proposed in Ref. Sugimoto:2025xva .
3 Numerical results
@@
3.1 Chiral condensate and number density
@@
Figure 2 plots $\langle{\bar{\chi}}\chi\rangle$ and $\langle n\rangle$, together with the diquark condensate $\langle\chi\chi\rangle$, as a function of $\mu$ at $m=1.0$ on a $1024^{4}$ lattice with the choice of $D=55$. We find that the qualitative behaviors of $\langle{\bar{\chi}}\chi\rangle$ and $\langle n\rangle$ are similar to those for $m\neq 0$ given by the MF analysis in Ref. Nishida:2003uj : $\langle{\bar{\chi}}\chi\rangle$ stays constant up to $\mu=\mu_{c}^{\rm low}$ and monotonically decreases up to the vanishing point of $\mu=\mu_{c}^{\rm up}$, where $\mu_{c}^{\rm low}\leq\mu\leq\mu_{c}^{\rm up}$ is the region of $\langle\chi\chi\rangle\neq 0$; $\langle n\rangle/2$ starts to increase from zero at $\mu=\mu_{c}^{\rm low}$ and saturates to $\langle n\rangle/2=1$ at $\mu\geq\mu_{c}^{\rm up}$. Note that $\langle n\rangle>2$ is not allowed due to the Pauli exclusion principle.
@@
3.2 Diquark condensate
@@
where the coefficient $b_{2}$ gives the diquark condensate. The range of $\lambda$ was chosen depending on $\mu$: for $\mu=1.095$, $\lambda\in[0,0.030]$ with $\Delta\lambda=0.002$, for $1.097\leq\mu\leq 1.11$, $\lambda\in[0,0.045]$ with $\Delta\lambda=0.005$, and for other $\mu$, $\lambda\in[0,0.040]$ with $\Delta\lambda=0.005$. We find that the $\mu$ dependence of the diquark condensate in Fig. 2 is qualitatively described by the MF analysis in Ref. Nishida:2003uj . The region of $\langle\chi\chi\rangle\neq 0$ shows rough consistency with the MF prediction, which is denoted by the gray band from $\mu_{c}^{\rm low}=1.0913$ to $\mu_{c}^{\rm up}=1.1944$.
Now let us determine the critical exponents $\beta_{m}$ and $\delta$. Figure 3 shows the $\mu$ dependence of $\langle\chi\chi\rangle$ with $D=55$ near the critical point $\mu_{c}^{\rm low}$. A fit of $\langle\chi\chi\rangle$ in the region of $\mu\in[1.10,1.12]$ assuming $\langle\chi\chi\rangle=A(\mu-\mu_{c}^{\rm low})^{\beta_{m}}$ with $A$, $\mu_{c}^{\rm low}$ and $\beta_{m}$ the free parameters yields $A=4.7(4)$, $\mu_{c}^{\rm low}=1.0950(7)$ and $\beta_{m}=0.514(27)$. The value of the critical $\mu$ is close to the MF prediction of $\mu_{c}^{\rm low}=1.0913$ and $\beta_{m}$ shows consistency with the MF prediction of $\beta_{m}=0.5$. The lower limit of the fitting range is chosen to avoid the immediate vicinity of the critical point. Below $\mu=1.10$ non-analytic behaviors of $f(m,\mu,\lambda)$ in terms of $\lambda$ becomes manifest. The robustness of the fit results against the change of the upper limit of the fitting range is indicated by the fact that the data in the region of $1.12<\mu\leq 1.13$ is on the fit curve. We also estimate another critical exponent $\delta$ from the $\lambda$ dependence of the free energy at $\mu_{c}^{\rm low}=1.095$ plotted in Fig. 4. The solid curve denotes the fit results of the form $f(m,\mu,\lambda=0)+b_{0}\lambda^{1+1/\delta}$ choosing the fitting range of $\lambda\in[0,0.03]$. The free parameters are determined as $b_{0}=1.51(6)$ and $\delta=2.44(6)$. The value of $\delta$ is slightly smaller than the MF prediction of $\delta=3$. The discrepancy may be attributed to the difficulty in the precise determination of $\mu_{c}^{\rm low}$.
@@
Figure 4: $\lambda$ dependence of $\langle\chi\chi\rangle$ at $(m,\mu)=(1.0,1.095)$ on a $1024^{4}$ lattice with $D=55$.
4 Summary and outlook

### v2

# Phase structure of (3+1)-dimensional dense two-color QCD at $T=0$ in the strong coupling limit with the tensor renormalization group

@@ Abstract
We investigate the phase structure of the (3+1)-dimensional strong coupling two-color QCD at zero temperature ($T=0$) with finite chemical potential using the tensor renormalization group method. The chiral and diquark condensates and the quark number density are evaluated as a function of the chemical potential. We further determine the critical exponents associated with the diquark condensate, which suggest consistency with the predictions of mean-field theory.
preprint: APS/123-QEDpreprint: UTHEP-811, UTCCS-P-169
I Introduction
@@
The complex action problem with finite chemical potential $\mu$ does not allow the standard Monte Carlo simulation to investigate the phase structure of QCD in the finite baryon density at $T=0$. On the other hand, the simulation of two-color QCD (QC2D) with $\mu\neq 0$ is free from the complex action problem, so that it provides us an opportunity to understand the physics of matter under the finite density condition. In the strong coupling limit, the analytical methods have been used to investigate the phase structure. In the 1980s, there were several mean field (MF) studies on chiral and diquark condensates in the strong coupling limit of the finite density QC2D with the Kogut-Susskind quark [1, 2, 3]. The results were later refined by Nishida, including the QCD case [4, 5]. On the other hand, at finite coupling the standard Monte Carlo approaches have been employed since the mid-1980s [6]. Although large-scale simulations are attainable in recent years, systematic studies changing the lattice sizes reveal temperature dependence of the phase structure [7, 8, 9]. It is manifest that larger lattice sizes are required toward zero temperature limit.
In this paper we investigate the phase structure in the strong coupling limit of the (3+1)-dimensional ((3+1)$d$) finite density QC2D at $T=0$ using the tensor renormalization group (TRG) method  [^1] In this paper, the “TRG method” or the “TRG approach” refers to not only the original numerical algorithm proposed by Levin and Nave [47] but also its extensions [48, 11, 49, 50, 51, 52, 37, 53, 23, 35, 36].. With the use of the Kogut-Susskind quark, we measure the chiral and diquark condensates and the quark number density as a function of $\mu$. There are two purposes in this work. First, we enlarge the lattice size up to $1024^{4}$ with the TRG method to investigate the phase structure at $T=0$, where the standard Monte Carlo approach is hardly accessible. It is useful to compare our results with the analytical ones obtained by the MF method and the $1/d_{s}$ expansion with $d_{s}$ the spatial dimension [4]. Second, this is an appropriate preparatory study before exploring (3+1)$d$ finite density QCD, whose complex action problem prohibits the standard Monte Carlo approaches. The TRG method may be qualified to investigate the phase structure of finite density QCD: This method, which is in principle free from the sign problem or the complex action problem, has been successfully applied to various types of models with these problems [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. So far there exist several TRG applications, particularly dealing with the (1+1)$d$ QCD-like theories [33, 34], the (1+1)$d$ QCD in the strong coupling limit [26], and an effective field theory of the QCD at finite density [23]. In Ref. [34], the bond-weighted TRG (BTRG) algorithm [35, 36] is employed to study the finite-density QC2D. Although the tensor network formulation of the (3+1)$d$ QC2D can be straightforwardly obtained by following Ref. [34], we need to resort a different TRG algorithm rather than the BTRG, which assumes the (1+1)$d$ tensor network representation. A practically useful algorithm for the (3+1)$d$ QC2D can be the anisotropic TRG algorithm [37, 38], which has been already applied to the (3+1)$d$ finite density Nambu-Jona-Lasinio (NJL) model [23]. Although computational techniques treating four-fermi interactions are in common between the NJL model and the strong coupling QC2D, the introduction of the color degree of freedom significantly increases the computational difficulty enlarging the size of the initial tensor by $2^{16}$. Since this difficulty is further enhanced drastically from QC2D to QCD we need to develop an efficient method to treat the large size of the initial tensor.
@@
This paper is organized as follows. In Sec. II, we define the action of the QC2D with finite chemical potential and give the tensor network representation. In Sec. III we measure the chiral and diquark condensates and the quark number density as a function of the chemical potential. The critical exponents are also determined. The results are compared with the analytical ones in Ref. [4]. Section IV is devoted to the summary and outlook.
II Formulation and numerical algorithm
II.1 Strong coupling QC2D
@@
We consider the finite density QC2D on a (3+1)$d$ lattice $\Lambda_{3+1}=\{(n_{1},n_{2},n_{3},n_{4})\ |n_{1,2,3}=1,\dots,N_{s}\,,n_{4}=1,\dots,N_{\tau}\}$ whose volume is $V=N_{s}^{3}\times N_{\tau}$ and temperature is defined as $T=1/N_{\tau}$. The lattice spacing $a$ is set to $a=1$ unless necessary. The action consists of the gluonic part and the fermionic one. For the former, the simple plaquette action is given by
@@
with $m$ the quark mass. $\chi=(\chi^{1},\chi^{2},\dots,\chi^{N_{c}})$ is a $N_{c}$ component Grassmann variable, and $\eta_{\nu}(n)$ is the staggered sign function defined by $\eta_{\nu}(n)=(-1)^{n_{1}+\cdots+n_{\nu-1}}$ with $\eta_{1}(n)=1$.
It is worthwhile to describe the symmetries in this theory. Let us start with the $m=\mu=0$ case. Equation (2), which corresponds to four-flavor quarks in the continuum limit, is invariant under the following continuous vector and axial vector transformation:
$\chi(n)\rightarrow e^{i\alpha}\chi(n),\;\;\bar{\chi}(n)\rightarrow\bar{\chi}(n)e^{-i\alpha}$ (3)
$\chi(n)\rightarrow e^{i\alpha\epsilon(n)}\chi(n),\;\;\bar{\chi}(n)\rightarrow\bar{\chi}(n)e^{i\alpha\epsilon(n)}$ (4)
[…]
@@
where we add an explicit U(1)V symmetry breaking term $S_{D}$ to detect the diquark condensation [39, 42]:
$S_{D}=-\frac{\lambda}{2}\{\chi^{t}\tau_{2}\chi+\bar{\chi}\tau_{2}{\bar{\chi}}^{t}\}$ (6)
@@
II.2 Grassman tensor network representation
In order to apply tensor network methods to this system, it is necessary to rewrite the partition function $Z$ in Eq. (5) as a product of local tensors. Since the theory involves fermionic degrees of freedom, we must properly incorporate their anticommuting nature. To this end, we employ a Grassmann tensor network formulation [43], in which the local tensors are defined using auxiliary Grassmann variables.
For clarity, let us consider the case $\lambda=0$, which simplifies the discussion. The extension to $\lambda\neq 0$ is straightforward. To obtain the Grassmann tensor network representation of the partition function in Eq. (5), we introduce two-component auxiliary Grassmann variables $\zeta_{\nu}$ and $\xi_{\nu}$ $(\nu=1,2,3,4)$ [43]. As a basic tool, for any $N$-component Grassmann variable $\theta_{i}$, one can use the identity
@@ II.2 Grassman tensor network representation
In the above, we have introduced the fermion occupation numbers $i^{\prime 1}_{\nu},i^{\prime 2}_{\nu},j^{\prime 2}_{\nu},j^{\prime 1}_{\nu},j^{1}_{\nu},j^{2}_{\nu},i^{1}_{\nu},i^{2}_{\nu}\in\{0,1\}$ $(\nu=1,2,3,4)$, which arise from the Taylor expansion of Eq. (10), and the auxiliary integer variables $a_{\nu},b_{\nu},c_{\nu},d_{\nu}\in\{1,2\}$ to sum out the color indices of the staggered fields $\chi$ connected to the link variables $U_{\nu}$. The summation $\sum_{\{i,j\}}$ denotes $\left(\prod_{\nu}\sum_{i^{\prime 1}_{\nu},i^{\prime 2}_{\nu},j^{\prime 1}_{\nu},j^{\prime 2}_{\nu},j^{1}_{\nu},j^{2}_{\nu},i^{1}_{\nu},i^{2}_{\nu}}\right)$. Here we omit the explicit spacetime dependence except for $\eta_{\nu}(n)$. This implies that there are eight types of initial tensors, depending on the staggered sign [23].
@@
where $F^{a_{\nu}b_{\nu}c_{\nu}d_{\nu}}_{i^{\prime 2}_{\nu}i^{\prime 1}_{\nu}j^{\prime 2}_{\nu}j^{\prime 1}_{\nu}}$ indicates the results of the integration using Weingarten calculus [44]. After integrating out the gauge fields, all that remains is to integrate out $\chi$ and $\bar{\chi}$. A straightforward calculation yields the final form of the Grassmann tensor $\mathcal{T}$:
@@
The tensor $\tilde{F}$ is a redefined version of $F$ that incorporates $a,b,c,d$ dependence of the sign factors arising from the reordering of $\chi^{2},\chi^{1},\bar{\chi}^{2},\bar{\chi}^{1}$ in the integration procedure. The tensor $R$ likewise includes sign factors originating from the rearrangement of the auxiliary Grassmann fields $\zeta,\xi,\bar{\zeta},\bar{\xi}$ to the form of Eq. (14). The explicit definitions of $F$, $\tilde{F}$, and $R$ are provided in the Appendix. Using these definitions, the partition function $Z$ can be expressed as a Grassmann tensor network [18],
@@
To evaluate Eq. (16), we have employed the Grassmann-anisotropic TRG method [37, 23] with the bond dimension $D$ incorporating the multi-GPU parallelization technique proposed in Ref. [45].
III Numerical results
@@
III.1 Chiral condensate and number density
@@
Figure 2 plots $\langle{\bar{\chi}}\chi\rangle$ and $\langle n\rangle$, together with the diquark condensate $\langle\chi\chi\rangle$, as a function of $\mu$ at $m=1.0$ on a $1024^{4}$ lattice with the choice of $D=55$. We find that the qualitative behaviors of $\langle{\bar{\chi}}\chi\rangle$ and $\langle n\rangle$ are similar to those for $m\neq 0$ given by the MF analysis in Ref. [4]: $\langle{\bar{\chi}}\chi\rangle$ stays constant up to $\mu=\mu_{c}^{\rm low}$ and monotonically decreases up to the vanishing point of $\mu=\mu_{c}^{\rm up}$, where $\mu_{c}^{\rm low}\leq\mu\leq\mu_{c}^{\rm up}$ is the region of $\langle\chi\chi\rangle\neq 0$; $\langle n\rangle/2$ starts to increase from zero at $\mu=\mu_{c}^{\rm low}$ and saturates to $\langle n\rangle/2=1$ at $\mu\geq\mu_{c}^{\rm up}$. Note that $\langle n\rangle>2$ is not allowed due to the Pauli exclusion principle. In Fig. 3 we plot the results for $\langle{\bar{\chi}}\chi\rangle$ and $\langle n\rangle$ changing the volume. It is clear that the $1024^{4}$ lattice size is large enough to obtain the results in the thermodynamic limit at the zero temperature.
@@
Figure 3: Volume dependence of $\langle{\bar{\chi}}\chi\rangle$ and $\langle n\rangle/2$ at $(m,\lambda)=(1.0,0)$ with $D=55$.
III.2 Diquark condensate
@@
where the coefficient $b_{2}$ gives the diquark condensate. The range of $\lambda$ was chosen depending on $\mu$: for $\mu=1.095$, $\lambda\in[0,0.030]$ with $\Delta\lambda=0.002$, for $1.097\leq\mu\leq 1.11$, $\lambda\in[0,0.045]$ with $\Delta\lambda=0.005$, and for other $\mu$, $\lambda\in[0,0.040]$ with $\Delta\lambda=0.005$. We find that the $\mu$ dependence of the diquark condensate in Fig. 2 is qualitatively described by the MF analysis in Ref. [4]. The region of $\langle\chi\chi\rangle\neq 0$ shows rough consistency with the MF prediction, which is denoted by the gray band from $\mu_{c}^{\rm low}=1.0913$ to $\mu_{c}^{\rm up}=1.1944$. Figure 4 shows the volume dependence of the diquark condensate. We hardly observe the volume dependence for the results on $V\geq 2^{16}$. As in the case of $\langle{\bar{\chi}}\chi\rangle$ and $\langle n\rangle$, the diquark condensate in the thermodynamic limit at zero temperature is obtained on $V=1024^{4}$.
[…]
Figure 4: Volume dependence of $\langle\chi\chi\rangle$ at $(m,\lambda)=(1.0,0)$ with $D=55$.
@@
Figure 6: $\lambda$ dependence of $\langle\chi\chi\rangle$ at $m=1.0$ on a $1024^{4}$ lattice with $D=55$. Circle, triangle and square symbols denote the results at $\mu=1.0950$, 1.0957 and $1.0943$, respectively.
IV Summary and outlook